# The gauge coupling evolutions of an ${\rm SU}(8)$ theory with the maximally symmetry breaking pattern

### v1

@@ 2.4 Decompositions of the ${\rm SU}(8)$ Higgs fields
We decompose the Higgs fields in Eq. (6) into components that can be responsible for the sequential symmetry breaking pattern in Eq. (1). For Higgs fields of $\mathbf{\overline{8_{H}}}_{\,,\omega}$, they read
@@
Table 5: The Higgs fields and their symmetry breaking directions in the ${\rm SU}(8)$ theory. The ✓and ✗represent possible and impossible symmetry breaking directions for a given Higgs field. The flavor indices for the Higgs fields that develop VEVs at different stages are also specified in the parentheses.
All components that are likely to develop VEVs for the sequential symmetry breaking stages are framed with boxes, and other ${\rm SU}(3)_{c}\otimes{\rm U}(1)_{\rm EM}$-exact components are neglected. Accordingly, we mark the allowed/disallowed symmetry breaking directions for all Higgs fields in Tab. 5. Schematically, we assign the Higgs VEVs according to the symmetry breaking pattern in Eq. (1) as follows
@@ 2.5 A summary of the vectorlike fermions of the SWW symmetry breaking pattern
Here, we summarize all vectorlike fermions that acquire masses above the EWSB scale in Tab. 6, with the conventions of flavor indices displayed explicitly.
@@ 3 The gauge sector of the ${\rm SU}(8)$ theory and the flavor non-universality
In this section, we describe the gauge sector according to the SWW symmetry breaking pattern in Eq. (1). The gauge couplings at each symmetry breaking stage will be defined for the derivation of RGEs. Due to the intrinsic symmetry properties, the flavor non-universality of three-generational SM fermions in the $\mathcal{G}_{441}$ and the $\mathcal{G}_{341}$ theories will be explicitly displayed. Some of the earlier and recent references addressing the flavor non-universality scenario with product gauge groups in the UV setup include [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
@@ 3.1.1 Covariant derivatives and gauge boson masses
The ${\rm SU}(4)_{s}$ generators of $T_{{\rm SU}(4)}^{\bar{A}}$ are normalized such that ${\rm Tr}\Big{(}T_{{\rm SU}(4)}^{\bar{A}}T_{{\rm SU}(4)}^{\bar{B}}\Big{)}=\frac{1}{2}\delta^{\bar{A}\bar{B}}$, with $\bar{A}\,,\bar{B}=1\,,...\,,15$. Specifically, we split the ${\rm SU}(4)_{s}$ adjoint indices into $\bar{A}=\underbrace{1\,,...\,,8}_{A}\,,\underbrace{9\,,...\,,15}_{\hat{A}}$. All ${\rm SU}(4)_{s}$ gluons of $G_{\mu}^{\bar{A}}$ are decomposed into eight massless ${\rm SU}(3)_{c}$ adjoint gluons, three massive ${\rm SU}(3)_{c}$ fundamental and anti-fundamental vector bosons of $(\tilde{C}_{\mu}^{a}\,,\tilde{C}_{\mu\,,a})$, and one massive ${\rm SU}(3)_{c}$-singlet vector boson of $M_{\mu}$. It is useful to define a strong mixing angle of
@@ 3.1.2 Gauge couplings of the $\mathcal{G}_{441}$ fermions
and we tabulate the vectorial and axial couplings of $(g_{f}^{V\,\mathcal{M}}\,,g_{f}^{A\,\mathcal{M}})$ in Tab. 7. The flavor non-universality of the neutral currents mediated by $M_{\mu}$ among three generational SM quarks and leptons are displayed explicitly.
@@ 3.2.2 Gauge couplings of the $\mathcal{G}_{341}$ fermions
$\supset$ $\frac{g_{X_{2}}}{s_{\vartheta_{G}}c_{\vartheta_{G}}}\Big{[}(\frac{1}{6}+\frac{1}{3}s_{\vartheta_{G}}^{2})(\overline{\check{\mathfrak{n}}_{R}^{\prime\prime\prime}}\gamma^{\mu}\check{\mathfrak{n}}_{R}^{\prime\prime\prime}+\overline{\mathfrak{n}_{R}^{\prime\prime\prime\prime}}\gamma^{\mu}\mathfrak{n}_{R}^{\prime\prime\prime\prime}+\overline{\mathfrak{e}_{R}^{\prime\prime\prime\prime}}\gamma^{\mu}\mathfrak{e}_{R}^{\prime\prime\prime\prime})$
@@
Table 8: The couplings of the flavor-conserving neutral currents mediated by $Z_{\mu}^{\prime\prime}$ in the $V-A$ basis.
@@ 3.3.1 Covariant derivatives and gauge boson masses
$\alpha_{2W}^{-1}(v_{331})=\alpha_{3W}^{-1}(v_{331})\,,\leavevmode\nobreak\ \alpha_{Y}^{-1}(v_{331})=\frac{1}{3}\alpha_{3W}^{-1}(v_{331})+\alpha_{X_{2}}^{-1}(v_{331})\,,$ (153)
$\frac{1}{3}\alpha_{3W}^{-1}=\alpha_{Y}^{-1}s_{\vartheta_{S}}^{2}\,,\leavevmode\nobreak\ \alpha_{X_{2}}^{-1}=\alpha_{Y}^{-1}c_{\vartheta_{S}}^{2}\,.$
@@
from the gauge coupling matching in Eq. (153). The tree-level masses for five gauge bosons at this stage read
@@ 3.4.2 Fermion $\mathcal{G}_{\rm SM}$ gauge couplings
4 The gauge coupling unification in the ${\rm SU}(8)$ theory
@@
4.1 RGEs of the minimal setup
@@
$(\mathbf{4}\,,\mathbf{\overline{4}}\,,+\frac{1}{2})_{\mathbf{H}}\oplus(\mathbf{\overline{4}}\,,\mathbf{4}\,,-\frac{1}{2})_{\mathbf{H}}\oplus(\mathbf{6}\,,\mathbf{6}\,,0)_{\mathbf{H}}\subset\mathbf{70_{H}}\,.$ (184)
@@
Figure 1: The RGEs of the minimal ${\rm SU}(8)$ setup. The dashed lines and the solid lines represent the one-loop and two-loop RGEs, respectively.
With the one- and two-loop $\beta$ coefficients defined in Eqs. (188), (194), (200), and (4.1), we plot the RGEs of the minimal ${\rm SU}(8)$ setup in Fig. 1. Three intermediate symmetry breaking scales follow from the benchmark point in Eq. (2). The discontinuities of the ${\rm U}(1)$ gauge couplings at three intermediate scales follow from their definitions in Eqs. (3.1.1), (3.2.1), (153), respectively. The RG evolutions within $10^{3}\,{\rm GeV}\lesssim\mu\lesssim 10^{13}\,{\rm GeV}$ are zoomed out in order to highlight the behaviors in three intermediate symmetry breaking scales. The current plot indicates $\alpha_{4S}^{-1}(\mu)\sim\alpha_{4W}^{-1}(\mu)\simeq 40$ and $\alpha_{X_{0}}^{-1}(\mu)\simeq 20$, for $v_{441}\lesssim\mu\lesssim 10^{18}\,{\rm GeV}$. Furthermore, the one-loop $\beta$ coefficients of $b_{{\rm SU}(4)_{s}}^{(1)}=b_{{\rm SU}(4)_{W}}^{(1)}<b_{{\rm U}(1)_{X_{0}}}^{(1)}$ from Eq. (188) guarantee that two non-Abelian ${\rm SU}(4)_{S\,,W}$ gauge couplings can never unify with the Abelian ${\rm U}(1)_{X_{0}}$ gauge coupling in the minimal setup.
@@ 4.2 The gauge coupling unification with additional adjoint Higgs fields
Accordingly, we find the following results
@@ 5.1 The main results
We study the RGEs of the gauge couplings in an ${\rm SU}(8)$ theory through its SWW symmetry breaking pattern in Eq. (1). Three intermediate symmetry breaking scales follow from the previous Ref. [10], which were obtained such that three generational SM quark/lepton masses and the CKM mixing pattern are consistently predicted. The gauge couplings at each symmetry breaking stage are defined in Sec. 3. Together with the symmetry breaking scales determined in Eq. (2), we obtain the RG evolution of three gauge couplings in Fig. 1 for the minimal setup. The gauge coupling unification cannot be achieved even when we include the one-loop threshold effects into account. The only possible solution within the field theory is to include addition $150-200$ adjoint massless Higgs fields of $(\mathbf{15}\,,\mathbf{1}\,,0)_{\mathbf{H}}\oplus(\mathbf{1}\,,\mathbf{15}\,,0)_{\mathbf{H}}$ within the range of $v_{441}\lesssim\mu\lesssim v_{U}$.

### v2

@@ 2.4 Decompositions of the ${\rm SU}(8)$ Higgs fields
We decompose the Higgs fields in Eq. (7) into components that can be responsible for the sequential SWW symmetry breaking pattern in Eq. (1). For Higgs fields of $\mathbf{\overline{8_{H}}}_{\,,\omega}$, they read
@@
Table 5: The Higgs fields and their SWW symmetry breaking directions in the ${\rm SU}(8)$ theory. The ✓and ✗represent possible and impossible symmetry breaking directions for a given Higgs field. The flavor indices for the Higgs fields that develop VEVs at different stages are also specified in the parentheses.
All components that are likely to develop VEVs for the sequential symmetry breaking stages are framed by boxes, and other ${\rm SU}(3)_{c}\otimes{\rm U}(1)_{\rm EM}$-exact components are neglected. Accordingly, we mark the allowed/disallowed symmetry breaking directions for all Higgs fields in Tab. 5. Schematically, we assign the Higgs VEVs according to the symmetry breaking pattern in Eq. (1) as follows
@@ 2.5 A summary of the vectorlike fermions of the SWW symmetry breaking pattern
Here, we summarize all vectorlike fermions that acquire masses above the EWSB scale in Tab. 6, where the conventions of flavor indices aligned with our previous derivations in Refs. [8, 10]. In the language of Georgi’s decomposition and count rule [5], the $(\mathbf{5_{F}}\,,\mathbf{\overline{5_{F}}})$ vectorlike pairs appear in the form of $(\mathfrak{D}\,,\mathfrak{e}^{\prime\prime}\,,\mathfrak{n}^{\prime\prime})$ at the first stage, and so on. At the second symmetry breaking stage, there is also a $(\mathbf{10_{F}}\,,\mathbf{\overline{10_{F}}})$ vectorlike pair, which appear in the form of $(\mathfrak{u}\,,\mathfrak{d}\,,\mathfrak{U}\,,\mathfrak{E})$. The numbers of vectorlike pairs that become massive at each symmetry breaking stage have been precisely determined by gauge invariance and the anomaly-free conditions [8, 10].
@@ 3 The gauge sector of the ${\rm SU}(8)$ theory and the flavor non-universality
In this section, we describe the gauge sector according to the SWW symmetry breaking pattern in Eq. (1). The gauge couplings at each symmetry breaking stage will be defined for the derivation of RGEs. Due to the intrinsic symmetry properties, the flavor non-universality of three-generational SM fermions in the $\mathcal{G}_{441}$ and the $\mathcal{G}_{341}$ theories will be explicitly displayed. Some of the earlier and recent references addressing the flavor non-universality scenario with product gauge groups in the UV setup include Refs. [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36].
@@ 3.1.1 Covariant derivatives and gauge boson masses
The ${\rm SU}(4)_{s}$ generators of $T_{{\rm SU}(4)}^{\bar{A}}$ are normalized such that ${\rm Tr}\Big{(}T_{{\rm SU}(4)}^{\bar{A}}T_{{\rm SU}(4)}^{\bar{B}}\Big{)}=\frac{1}{2}\delta^{\bar{A}\bar{B}}$, with $\bar{A}\,,\bar{B}=1\,,...\,,15$. Specifically, we split the ${\rm SU}(4)_{s}$ adjoint indices into $\bar{A}=\underbrace{1\,,...\,,8}_{A}\,,\underbrace{9\,,...\,,15}_{\hat{A}}$. All ${\rm SU}(4)_{s}$ gluons of $G_{\mu}^{\bar{A}}$ are decomposed into
$\underbrace{(\mathbf{15}\,,\mathbf{1}\,,0)_{\mathbf{V}}}_{G_{\mu}^{\bar{A}}}$ $=$ $\underbrace{(\mathbf{8}\,,\mathbf{1}\,,0)_{\mathbf{V}}}_{G_{\mu}^{A}}\oplus\underbrace{(\mathbf{3}\,,\mathbf{1}\,,-\frac{1}{3})_{\mathbf{V}}}_{\tilde{C}_{\mu}^{a}}\oplus\underbrace{(\mathbf{\overline{3}}\,,\mathbf{1}\,,+\frac{1}{3})_{\mathbf{V}}}_{\tilde{C}_{\mu\,,a}}\oplus\underbrace{(\mathbf{1}\,,\mathbf{1}\,,0)_{\mathbf{V}}}_{M_{\mu}}\,.$ (22)
It is useful to define a strong mixing angle of
@@ 3.1.2 Gauge couplings of the $\mathcal{G}_{441}$ fermions
and we tabulate the vectorial and axial couplings of $(g_{f}^{V\,\mathcal{M}}\,,g_{f}^{A\,\mathcal{M}})$ in Tab. 7. The flavor non-universality of the neutral currents mediated by $M_{\mu}$ among three generational SM quarks and leptons are displayed explicitly. Three generational SM left-handed neutrinos exhibit flavor universal coupling to the $M_{\mu}$, since they all originate from the identical irrep of $\mathbf{\overline{8_{F}}}$. This will always be true in their couplings to the neutral gauge bosons in the sequential symmetry breaking stages.
@@ 3.2.2 Gauge couplings of the $\mathcal{G}_{341}$ fermions
Table 8: The couplings of the neutral currents mediated by $Z_{\mu}^{\prime\prime}$ in the $V-A$ basis.
@@ 3.3.1 Covariant derivatives and gauge boson masses
$\alpha_{2W}^{-1}(v_{331})=\alpha_{3W}^{-1}(v_{331})\,,\leavevmode\nobreak\ \alpha_{Y}^{-1}(v_{331})=\frac{1}{3}\alpha_{3W}^{-1}(v_{331})+\alpha_{X_{2}}^{-1}(v_{331})\,,$
$\frac{1}{3}\alpha_{3W}^{-1}=\alpha_{Y}^{-1}s_{\vartheta_{S}}^{2}\,,\leavevmode\nobreak\ \alpha_{X_{2}}^{-1}=\alpha_{Y}^{-1}c_{\vartheta_{S}}^{2}\,.$ (154)
@@
from the gauge coupling matching in Eq. (3.3.1). The tree-level masses for five gauge bosons at this stage read
@@ 3.4.2 Fermion $\mathcal{G}_{\rm SM}$ gauge couplings
4 The gauge coupling evolutions in the ${\rm SU}(8)$ theory
@@
4.1 RGEs of the minimal setup with the survival hypothesis
@@
$(\mathbf{1}\,,\mathbf{1}\,,-1)_{\mathbf{H}}^{\prime\prime}\oplus(\mathbf{1}\,,\mathbf{1}\,,+1)_{\mathbf{H}}^{\prime\prime\prime\prime}\oplus(\mathbf{4}\,,\mathbf{\overline{4}}\,,+\frac{1}{2})_{\mathbf{H}}\oplus(\mathbf{\overline{4}}\,,\mathbf{4}\,,-\frac{1}{2})_{\mathbf{H}}\oplus(\mathbf{6}\,,\mathbf{6}\,,0)_{\mathbf{H}}\subset\mathbf{70_{H}}\,.$ (185)
@@
When the RGEs evolve down to $v_{441}$, the gauge couplings should match according to Eq. (3.1.1).
@@
When the RGEs evolve down to $v_{341}$, the gauge couplings should match according to Eq. (3.2.1).
@@
When the RGEs evolve down to $v_{331}$, the gauge couplings should match according to Eq. (3.3.1).
@@
Figure 1: The RGEs of the minimal ${\rm SU}(8)$ setup according to the survival hypothesis. The dashed lines and the solid lines represent the one-loop and two-loop RGEs, respectively.
With the one- and two-loop $\beta$ coefficients defined in Eqs. (189), (195), (201), and (4.1), we plot the RGEs of the minimal ${\rm SU}(8)$ setup in Fig. 1. Three intermediate symmetry breaking scales follow from the benchmark point in Eq. (2). The discontinuities of the ${\rm U}(1)$ gauge couplings at three intermediate scales follow from their definitions in Eqs. (3.1.1), (3.2.1), (3.3.1), respectively. The RG evolutions within $10^{3}\,{\rm GeV}\lesssim\mu\lesssim 10^{13}\,{\rm GeV}$ are zoomed out in order to highlight the behaviors in three intermediate symmetry breaking scales. The current plot indicates $\alpha_{4S}^{-1}(\mu)\sim\alpha_{4W}^{-1}(\mu)\simeq 40$ and $\alpha_{X_{0}}^{-1}(\mu)\simeq 20$ for $\mu\simeq 5\times 10^{17}\,{\rm GeV}$. Furthermore, the one-loop $\beta$ coefficients of $b_{{\rm SU}(4)_{s}}^{(1)}=b_{{\rm SU}(4)_{W}}^{(1)}<b_{{\rm U}(1)_{X_{0}}}^{(1)}$ from Eq. (189) guarantee that two non-Abelian ${\rm SU}(4)_{S\,,W}$ gauge couplings can never unify with the Abelian ${\rm U}(1)_{X_{0}}$ gauge coupling in the minimal setup.
@@ 4.2 The gauge coupling unification with additional adjoint Higgs fields
Therefore, we find the following results
@@
4.3 RGEs of the minimal setup with the lighter Higgs fields assumptions
The derivations in Sec. 4.1 relies on the minimal set of massless Higgs fields that can develop the VEVs according to Eqs. (18). Next, we consider the RGEs of the minimal setup by assuming more lighter Higgs fields relative to the survival hypothesis. In this scenario, the Higgs fields from the $\mathbf{\overline{8_{H}}}_{\,,\omega}\oplus\mathbf{\overline{28_{H}}}_{\,,\dot{\omega}}$ remain massless until certain components obtained VEVs according to Eqs. (18). Meanwhile, the massless fermions at each symmetry breaking stage have been fixed according to the anomaly-free conditions [8, 10].
Between the $v_{441}\leq\mu\leq v_{U}$, we consider the identical massless Higgs fields in Eq. (4.1). Between the $v_{341}\leq\mu\leq v_{441}$, we assume the massless Higgs fields of
$(\mathbf{\overline{3}}\,,\mathbf{1}\,,+\frac{1}{3})_{\mathbf{H}\,,3\,,{\rm V}\,,{\rm VI}}\oplus(\mathbf{1}\,,\mathbf{\overline{4}}\,,-\frac{1}{4})_{\mathbf{H}\,,3\,,{\rm V}\,,{\rm VI}}\subset\mathbf{\overline{8_{H}}}_{\,,3\,,{\rm V}\,,{\rm VI}}\,,$ (216)
$(\mathbf{3}\,,\mathbf{1}\,,+\frac{2}{3})_{\mathbf{H}\,,\dot{\omega}}\oplus(\mathbf{\overline{3}}\,,\mathbf{1}\,,+\frac{1}{3})_{\mathbf{H}\,,\dot{\omega}}\oplus(\mathbf{1}\,,\mathbf{6}\,,-\frac{1}{2})_{\mathbf{H}\,,\dot{\omega}}$
$\oplus$ $(\mathbf{\overline{3}}\,,\mathbf{\overline{4}}\,,+\frac{1}{12})_{\mathbf{H}\,,\dot{\omega}}\oplus(\mathbf{1}\,,\mathbf{\overline{4}}\,,-\frac{1}{4})_{\mathbf{H}\,,\dot{\omega}}\subset\mathbf{\overline{28_{H}}}_{\,,\dot{\omega}}\,,\quad\dot{\omega}=(\dot{1}\,,\dot{2}\,,\dot{\rm VII}\,,\dot{\rm VIII}\,,\dot{\rm IX})\,,$
$(\mathbf{1}\,,\mathbf{\overline{4}}\,,+\frac{3}{4})_{\mathbf{H}}^{\prime}\oplus(\mathbf{1}\,,\mathbf{4}\,,-\frac{3}{4})_{\mathbf{H}}^{\prime}\subset(\mathbf{4}\,,\mathbf{\overline{4}}\,,+\frac{1}{2})_{\mathbf{H}}\oplus(\mathbf{\overline{4}}\,,\mathbf{4}\,,-\frac{1}{2})_{\mathbf{H}}\subset\mathbf{70_{H}}\,,$
which leads to the $\mathcal{G}_{341}$ $\beta$ coefficients of
$(b_{{\rm SU}(3)_{c}}^{(1)}\,,b_{{\rm SU}(4)_{W}}^{(1)}\,,b_{{\rm U}(1)_{X_{1}}}^{(1)})=(+\frac{23}{6}\,,+\frac{1}{2}\,,+\frac{677}{36})\,,$ (220)
$b_{\mathcal{G}_{341}}^{(2)}=\left(\begin{array}[]{ccc}589/3&225/2&331/36\\ 60&509/2&45/4\\ 662/9&675/4&8893/216\\ \end{array}\right)\,.$
Between the $v_{331}\leq\mu\leq v_{341}$, we assume the massless Higgs fields of
$(\mathbf{\overline{3}}\,,\mathbf{1}\,,+\frac{1}{3})_{\mathbf{H}\,,3\,,{\rm VI}}\oplus(\mathbf{1}\,,\mathbf{\overline{3}}\,,-\frac{1}{3})_{\mathbf{H}\,,3\,,{\rm VI}}\subset\mathbf{\overline{8_{H}}}_{\,,3\,,{\rm VI}}\,,$ (221)
$\left[(\mathbf{3}\,,\mathbf{1}\,,+\frac{2}{3})_{\mathbf{H}\,,\dot{\omega}}\oplus(\mathbf{\overline{3}}\,,\mathbf{1}\,,+\frac{1}{3})_{\mathbf{H}\,,\dot{\omega}}\right]\oplus\left[(\mathbf{1}\,,\mathbf{\overline{3}}\,,-\frac{1}{3})_{\mathbf{H}\,,\dot{\omega}}^{\prime}\oplus(\mathbf{1}\,,\mathbf{3}\,,-\frac{2}{3})_{\mathbf{H}\,,\dot{\omega}}\right]$
$\oplus$ $\left[(\mathbf{\overline{3}}\,,\mathbf{\overline{3}}\,,0)_{\mathbf{H}\,,\dot{\omega}}\oplus(\mathbf{\overline{3}}\,,\mathbf{1}\,,+\frac{1}{3})_{\mathbf{H}\,,\dot{\omega}}\oplus(\mathbf{1}\,,\mathbf{\overline{3}}\,,-\frac{1}{3})_{\mathbf{H}\,,\dot{\omega}}\right]\subset\mathbf{\overline{28_{H}}}_{\,,\dot{\omega}}\,,\quad\dot{\omega}=(\dot{2}\,,\dot{\rm VIII}\,,\dot{\rm IX})\,,$
$(\mathbf{1}\,,\mathbf{\overline{3}}\,,+\frac{2}{3})_{\mathbf{H}}^{\prime\prime\prime}\subset(\mathbf{1}\,,\mathbf{\overline{4}}\,,+\frac{3}{4})_{\mathbf{H}}^{\prime}\subset(\mathbf{4}\,,\mathbf{\overline{4}}\,,+\frac{1}{2})_{\mathbf{H}}\subset\mathbf{70_{H}}\,.$
which leads to the $\mathcal{G}_{331}$ $\beta$ coefficients of
$(b_{{\rm SU}(3)_{c}}^{(1)}\,,b_{{\rm SU}(3)_{W}}^{(1)}\,,b_{{\rm U}(1)_{X_{2}}}^{(1)})=(-\frac{5}{3}\,,-\frac{3}{2}\,,+\frac{116}{9})\,,$ (225)
$b_{\mathcal{G}_{331}}^{(2)}=\left(\begin{array}[]{ccc}256/3&36&58/9\\ 36&89&22/3\\ 464/9&176/3&728/27\\ \end{array}\right)\,.$
Between the $v_{\rm EW}\leq\mu\leq v_{331}$, the massless Higgs field is identical to Eq. (202).
Figure 3: The RGEs of the minimal ${\rm SU}(8)$ setup with the lighter Higgs fields assumptions. The dashed lines and the solid lines represent the one-loop and two-loop RGEs, respectively.
Accordingly, we plot the RGEs with more lighter Higgs fields in Fig. 3. A very similar behavior can be found by comparing to the RGEs displayed in Fig. 1 with the survival hypothesis. The large discrepancy of $|\alpha_{X_{0}}^{-1}-(\alpha_{4S}^{-1}+\alpha_{4W}^{-1})/2|\sim 20$ beyond the $v_{441}$ can not be eliminated by considering the threshold effects as well. Without the inclusion of additional adjoint Higgs fields, the minimal ${\rm SU}(8)$ theory can not achieve the gauge coupling unification with two different assumptions of the Higgs fields masses.
@@ 5.1 The main results
We study the RGEs of the gauge couplings in an ${\rm SU}(8)$ theory through its SWW symmetry breaking pattern in Eq. (1). Three intermediate symmetry breaking scales follow from the previous Ref. [10], which were obtained such that three generational SM quark/lepton masses and the CKM mixing pattern are consistently predicted. The gauge couplings at each symmetry breaking stage are defined in Sec. 3. Since all three generational SM fermions are non-trivially embedded in the ${\rm SU}(8)$ theory, the flavor non-universality in terms of the neutral currents in the $\mathcal{G}_{441}$ and $\mathcal{G}_{341}$ effective theories are explicitly displayed in Tab. 7 and 8. Together with the symmetry breaking scales determined in Eq. (2), we obtain the RG evolution of three gauge couplings in Fig. 1 for the minimal setup. The gauge coupling unification cannot be achieved even when we include the one-loop threshold effects into account. The only possible solution within the field theory is to include addition $150-200$ adjoint massless Higgs fields of $(\mathbf{15}\,,\mathbf{1}\,,0)_{\mathbf{H}}\oplus(\mathbf{1}\,,\mathbf{15}\,,0)_{\mathbf{H}}$ within the range of $v_{441}\lesssim\mu\lesssim v_{U}$.